OpenStudy (ibbutibbu.):

Solve: 7x - 2(x + 1) = 6x + 14

OpenStudy (anonymous):

7x-2(x+1) = 6x+14 (original equation) 7x-2x-2 = 6x+14 (expand) 5x-2 = 6x+14 (combine like terms) -2 = x+14 (subtract 5x) x = -16 (subtract 14) Solution: x = -16

OpenStudy (luigi0210):

Start by distributing.

OpenStudy (anonymous):

7x - 2(x+1) = 6x + 14 7x - 2x - 2 = 6x + 14 5x - 2 = 6x + 14 5x = 6x + 16 -1x = 16 x = -16